ABC, Inc. manufactures three products: A, B, and C. Relevant data:

Unit Unit
Sales price Var Cost Sales in units
A 100 50 7,000
B 150 90 2,000
C 200 90 1,000
Fixed costs are $500,000.
Required:
a) What is the amount of profit at this sales volume?
b) What is the breakeven point (in sales dollars and units per product) at this sales mix?
c) What sales volume in dollars must be achieved to earn a profit of $200,000? How many units of Product C will be sold at this volume? (Assume constant sales mix)
d) Marketing department believes that a $10,000 promotional campaign for either B or C could result in a 20% one-time increase in
        sales for that product. Which product, if any, should be chosen? Explain.
e) Engineering department believes that acquiring new equipment for the Product A production line would reduct variable cost to $35/unit.
       This would allow the price to be cut to $90, resulting in sales of 8,000 units per year. Fixed costs would increase by $50,000/year.
      Sales of B and C would remain the same.
       Would this plan increase profits? Explain.
Answer #1

1)

Units

Sales price

Total Sales

Variable cost

Total VC

A

7000

100

700,000

50

350,000

B

2000

150

300,000

90

180,000

C

1000

200

200,000

90

90,000

1,200,000

620,000

Sales

1,200,000

Minus: VC

620,000

CM

580,000

Less: FC

500,000

Profit

80,000

CM = Sales - VC = 1,200,000 - 620,000 = 580,000

Profit = CM - FC = 580,000 - 5,00,000 = 80,000

2) Weighted average contribution per unit = CM/No of units = 580,000 / 10,000 = 58

Breakeven point in units = FC/Weighted average contribution per unit = 500,000 / 58 = 8620.59

No. of units of each product:

A = 8621 * 7/10 = 6035

B = 8621 * 2/10 = 1724

C = 8621 * 1/10 = 862

Breakeven point in $'s

A = 6035 * 100 =603,500

B = 1724 * 150 = 258,600

C = 862 * 200 = 172,400

Total = 603,500 + 258,600 + 172,400 = 1,034,500

3)

FC

500,000

Add: Desired Profit

200,000

700,000

PV ratio

48.33%

Sales needed to earn 20,000 profit

1448280

Total sales

1448280

Weighted average selling per unit

120

Units needed to earn 20,000 profit

1207

Weighted average selling per unit = Total sales / No of units = 1200,000 / 10,000 = 120

PV ratio = 580,000 = 1200,000 =48.33%

4)

Units

20% increase

Additional contribution

A

7000

1400

70,000

B

2000

2400

24,000

C

1000

1200

22,000

5)

VC

15

SP (100- 90)

-10

CM

5

Increase in sales

1,000

CM * Increase in sales

5,000

Additional FC

50,000

Profit/loss

-45,000